# Taming OOD Actions for Offline Reinforcement Learning: An Advantage-Based Approach

### v3

@@ 3.2 Advantage-Based Operator Analysis
Under this definition, an action $\bm{a}$ is considered advantageous if it leads to a next state with a higher expected value than the selective threshold defined by the $\kappa$-th quantile. A positive advantage indicates that the action is favored, while a negative advantage indicates that the action is penalized. The parameter $\kappa$ controls the level of conservatism: larger values lead to higher thresholds and encourage conservatism, while smaller values promote optimism by more readily rewarding unseen actions. Notably, all the components in Eq. (9) are learned solely from the offline dataset, providing a reliable basis for evaluating the quality of OOD actions.
@@
Offline RL algorithms based on the standard Bellman backup suffer from action distribution shift during training. This shift arises because the target values in Eq. (10) use actions sampled from the learned policy $\pi_{\theta}$, while the Q-function is trained only on actions sampled from the behavior policy that produced the dataset ($(\bm{s},\bm{a})\in\mathcal{D}$). Therefore, augmenting the standard Bellman operator with the advantage term $A(\bm{a^{\prime}}|\bm{s^{\prime}})$ can effectively mitigate estimation errors on OOD actions.
@@
The advantage-based Bellman operator defined in Eq. (10) is $\gamma$-contractive with respect to the $L_{\infty}$ norm, and any initial $Q$-function can converge to a unique fixed point by repeatedly applying this operator.
Therefore, we define the unique fixed point of Eq. (10) as $Q_{\pi_{\theta}}^{A}$, and denote by $Q_{\pi_{\theta}}$ the true Q-function of $\pi_{\theta}$. The relationship between $Q_{\pi_{\theta}}^{A}$ and $Q_{\pi_{\theta}}$ is characterized in the following proposition.
@@
$\|Q_{\pi_{\theta}}^{A}-Q_{\pi_{\theta}}\|_{\infty}\leq\frac{2\lambda R_{\rm max}}{(1-\gamma)^{2}}.$
Figure 2: Comparison of prior evaluation methods against our advantage-based evaluation. We visualize the Q-function and the policy for a fixed state. The thick blue solid line denotes the optimal value function. The solid line denotes the Q-function and the corresponding policy learned via advantage-based evaluation. The dashed line denotes the Q-function and policy learned via conservative evaluation. The dotted line denotes the Q-function and policy learned via standard Bellman backup evaluation.
Proposition 3.4 shows that the Q-function obtained by the advantage-based Bellman operator modifies the true Q-function derived from the standard Bellman operator within a bounded range of $2\lambda R_{\rm max}/(1-\gamma)^{2}$. We further illustrate the effectiveness of our advantage-based evaluation in Figure 2, where different line styles correspond to different evaluation methods. In the offline RL setting, due to the prohibition against interacting with the environment, directly applying standard Bellman backup (dotted line in Figure 2) results in an erroneous Q-function, which tends to overestimate the value of OOD actions and thus leads to an ineffective policy. Meanwhile, conservative evaluation (dashed line) indiscriminately penalizes all OOD actions, resulting in suppressed Q-values across these actions and limiting the learned policy to a sub-optimal policy near the support of the dataset. By contrast, our advantage-based evaluation (solid line) defines an advantage function that effectively modulates the Q-function obtained from Bellman backup, enabling the policy to discover optimal actions even beyond the support of the dataset. This phenomenon is further validated by our empirical results on the PointMaze tasks, as shown in Section 5.2.
@@ 4 Advantage-Based Diffusion Actor-Critic Algorithm
We now introduce our practical algorithm, Advantage-based Diffusion Actor-Critic (ADAC).
Diffusion Policy. We model our policy as the reverse process of a conditional diffusion model wang2022diffusion :
@@
where the terminal sample $a^{0}$ is used as the action for RL evaluation. During training, we sample $(\bm{s},\bm{a})$ pairs from the offline dataset $\mathcal{D}$ and construct noisy samples $\bm{a^{i}}=\sqrt{\bar{\alpha}_{i}}\bm{a}+\sqrt{1-\bar{\alpha}_{i}}\bm{\epsilon}$ (Eq. (3)), where $\alpha_{i}=1-\beta_{i}$, $\bar{\alpha}_{i}=\prod_{j=1}^{i}\alpha_{j}$, and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. Following DDPM ho2020denoising , we train the following noise prediction model $\bm{\epsilon}_{\theta}(\bm{a^{i}},\bm{s},i)$ to approximate the added noise, which determines the reverse process $p_{\theta}(\bm{a^{i-1}}\,|\,\bm{a^{i}},\bm{s})$:
$\mathcal{L}_{d}(\theta)=\mathbb{E}_{i\sim\mathcal{U},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}),(\bm{s},\bm{a})\sim\mathcal{D}}\left[\big{\|}\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{i}}\bm{a}+\sqrt{1-\bar{\alpha}_{i}}\bm{\epsilon},\bm{s},i)\big{\|}^{2}\right],$ (11)
where $\mathcal{U}$ denotes the uniform distribution over $\{1,\dots,T\}$. $\mathcal{L}_{d}(\theta)$ is a behavior cloning loss, and minimizing it enables the diffusion model to learn the behavior policy $\mu$. At inference time, an action $\bm{a^{0}}$ is generated by sampling $a^{T}\sim\mathcal{N}(\bm{0},\bm{I})$ and iteratively applying the learned reverse process.
@@
7:     Update $\mathcal{L}_{q}(\phi)$ according to Eq. (14) // Critic Update
8:     Update $\mathcal{L}_{p}(\theta)$ according to Eq. (15)// Actor Update
@@
Advantage Function. In our practical implementation, the value function is parameterized by a neural network with parameters $\varphi$ and trained by minimizing the following expectile regression loss:
$\mathcal{L}_{v}(\varphi)=\mathbb{E}_{(\bm{s},\bm{a},r,\bm{s^{\prime}})\sim\mathcal{D}}\big{[}L_{2}^{\tau}(r+\gamma V_{\varphi}(\bm{s^{\prime}})-V_{\varphi}(\bm{s}))\big{]}.$ (12)
@@
$\mathcal{L}_{p}(\psi)=\mathbb{E}_{(\bm{s},\bm{a},\bm{s^{\prime}})\sim\mathcal{D}}\big{[}\|P_{\psi}(\bm{s},\bm{a})-\bm{s^{\prime}}\|^{2}\big{]}.$ (13)
Therefore, the behavior policy $\mu$, the value function $V$, and the transition dynamics $P$ can be learned by minimizing Eq. (11), Eq. (12), and Eq. (13), respectively. The advantage function $A(\bm{a}\,|\,\bm{s})$ is then computed as defined in Eq. (9). All components are trained once using only offline data and remain fixed during the subsequent actor-critic updates, making this stage computationally inexpensive.
@@
$\mathcal{L}_{q}(\phi)=\mathbb{E}_{(\bm{s},\bm{a},\bm{s^{\prime}})\sim\mathcal{D},\bm{a^{\prime}}\sim\pi_{\theta}(\cdot\mid\bm{s^{\prime}})}\left[\left(r(\bm{s},\bm{a})+\gamma\big{(}Q_{\phi}(\bm{s^{\prime}},\bm{a^{\prime}})+\lambda A(\bm{a^{\prime}}\mid\bm{s^{\prime}})\big{)}-Q_{\phi}(\bm{s},\bm{a})\right)^{2}\right].$ (14)
@@
$\mathcal{L}_{p}(\theta)=\mathcal{L}_{d}(\theta)-\alpha\mathbb{E}_{\bm{s}\sim\mathcal{D},\bm{a}\sim\pi_{\theta}}\left[Q_{\phi}(\bm{s},\bm{a})\right].$ (15)
We summarize our implementation in Algorithm 1. A key innovation of our method lies in the use of $A(\bm{a}\,|\,\bm{s})$; without this component, our approach degenerates to the DQL method wang2022diffusion , which uses standard Bellman backup. In addition to achieving new SOTA, we compare with DQL on PointMaze in Section 5.2 to visualize how advantage modulation facilitates superior OOD action selection.
@@ 5 Experiments
In this section, we begin by evaluating our method on the widely recognized D4RL benchmark fu2020d4rl . We then design a dedicated experiment on PointMaze—another task within D4RL—to empirically and visually demonstrate ADAC’s ability to identify beneficial OOD actions. Finally, we perform an ablation study to dissect the contribution of key components in our method.
@@
Baseline. We consider a diverse array of baseline methods that exhibits strong results in each domain of tasks. For policy regularization-based method, we compare with the classic BC, TD3+BC fujimoto2021minimalist , BEARkumar2019stabilizing , BRACwu2019behavior , BCQfujimoto2019off , AWRpeng2019advantage , O-RLbrandfonbrener2021offline , and DQLwang2022diffusion . For pessimistic value function-based approach, we include CQLkumar2020conservative , IQLkostrikov2021offline , and REMagarwal2020optimistic . For model-based offline RL, we choose MoRelkidambi2020morel . For the classic online method, we include SAChaarnoja2018soft . For conditional sequence modeling approaches, we include DTchen2021decision , Diffuserjanner2022planning , and DDajay2022conditional . We report the performance of baseline methods either from the best results published in their respective papers or from wang2022diffusion .
@@
Gym Tasks BC TD3+BC CQL IQL MoRel DT Diffuser DD DQL ADAC (Ours) halfcheetah-medium 42.6 48.3 44.0 47.4 42.1 42.6 44.2 49.1 51.1 58.0$\pm$0.3 hopper-medium 52.9 59.3 58.5 66.3 95.4 67.6 58.5 79.3 90.5 93.5$\pm$4.2 walker2d-medium 75.3 83.7 72.5 78.3 77.8 74.0 79.7 82.5 87.0 87.6$\pm$2.0 halfcheetah-medium-replay 36.6 44.6 45.5 44.2 40.2 36.6 42.2 39.3 47.8 52.5$\pm$0.8 hopper-medium-replay 18.1 60.9 95.0 94.7 93.6 82.7 96.8 100.0 101.3 102.1$\pm$1.1 walker2d-medium-replay 26.0 81.8 77.2 73.9 49.8 66.6 61.2 75.0 95.5 96.0$\pm$1.6 halfcheetah-medium-expert 55.2 90.7 91.6 86.7 53.3 86.8 79.8 90.6 96.8 106.1$\pm$1.0 hopper-medium-expert 52.5 98.0 105.4 91.5 108.7 107.6 107.2 111.8 111.1 112.5$\pm$1.0 walker2d-medium-expert 107.5 110.1 108.8 109.6 95.6 108.1 108.4 108.8 110.1 112.3$\pm$0.9 Average 51.9 75.3 77.6 77.0 72.9 74.7 75.3 81.8 88.0 91.2 AntMaze Tasks BC TD3+BC CQL IQL BEAR DT BCQ O-RL DQL ADAC (Ours) antmaze-umaze 54.6 78.6 74.0 87.5 73.0 59.2 78.9 64.3 93.4 98.2$\pm$4.5 antmaze-umaze-diverse 45.6 71.4 84.0 62.2 61.0 53.0 55.0 60.7 66.2 76.0$\pm$9.9 antmaze-medium-play 0.0 10.6 61.2 71.2 0.0 0.0 0.0 0.3 76.6 86.5$\pm$9.8 antmaze-medium-diverse 0.0 3.0 53.7 70.0 8.0 0.0 0.0 0.0 78.6 88.7$\pm$10.2 antmaze-large-play 0.0 0.2 15.8 39.6 6.7 0.0 6.7 0.0 46.4 69.8$\pm$12.4 antmaze-large-diverse 0.0 0.0 14.9 47.5 2.2 0.0 2.2 0.0 56.6 64.6$\pm$12.7 Average 16.7 27.3 50.6 63.0 23.7 18.7 23.8 20.9 69.6 80.6 Adroit Tasks BC BRAC-v CQL IQL BEAR REM BCQ SAC DQL ADAC (Ours) pen-human 25.8 0.6 35.2 71.5 -1.0 5.4 68.9 4.3 72.8 74.4$\pm$18.6 pen-cloned 38.3 -2.5 27.2 37.3 26.5 -1.0 44.0 -0.8 57.3 80.5$\pm$14.3 Average 32.1 -1.0 31.2 54.4 12.8 2.2 56.5 1.8 65.1 77.5 Kitchen Tasks BC BRAC-v CQL IQL BEAR AWR BCQ SAC DQL ADAC (Ours) kitchen-complete 33.8 0.0 43.8 62.5 0.0 0.0 8.1 15.0 84.0 87.9$\pm$6.7 kitchen-partial 33.8 0.0 49.8 46.3 13.1 15.4 18.9 0.0 60.5 65.2$\pm$7.0 kitchen-mixed 47.5 0.0 51.0 51.0 47.2 10.6 8.1 2.5 62.6 68.3$\pm$5.8 Average 38.4 0.0 48.2 53.3 20.1 8.7 11.7 5.8 69.0 73.8
@@ 5.1 Benchmark Results
Our method is evaluated on four task domains, with results summarized in Table 1. We also provide domain-specific analysis to highlight key performance characteristics.
Results for Gym Tasks. While prior methods already achieve strong performance on Gym tasks, ADAC consistently delivers further improvements. Among all Gym tasks, halfcheetah is notably the most challenging, often yielding the lowest scores. However, ADAC achieves the largest performance boost on this task family, with an improvement of 10%. According to the protocol established in fu2020d4rl , a normalized score of 100 corresponds to an expert-level policy. Therefore, achieving an average normalized score exceeding 90 demonstrates the practical effectiveness of our method.
@@
Results for AntMaze Tasks. The AntMaze tasks are particularly challenging due to the sparse rewards and the prevalence of sub-optimal trajectories. Nevertheless, this is precisely where ADAC proves advantageous, as exploring OOD actions becomes essential for achieving high performance under such circumstances. Remarkably, we achieve over 15% improvements across all baseline methods. We visualize the learning process of the pre-SOTA method DQL and ADAC in Figure 3. As illustrated in Figure 3, ADAC consistently outperforms DQL in terms of both performance and stability across all three AntMaze task variants.
Results for Adroit and Kitchen Tasks. In the Adroit tasks, the narrow support of human demonstrations leads to substantial extrapolation error. While both DQL and ADAC incorporate policy regularization to address this issue, our advantage-based approach further mitigates it by effectively handling unseen actions beyond the demonstrations. As a result, we achieve a 20% performance improvement over existing baselines. Additionally, the Kitchen tasks require long-horizon value optimization, and we observe that ADAC also outperforms in this domain.

### v4

@@ 3.2 A New Advantage Function
Remark. The newly-defined advantage function is different from the more common one A(s,a):=Q(s,a)V(s)A(s,a):=Q(s,a)-V(s) which employs both the Q-function and V-function. However, relying on Q-function can typically cause over-estimation. In the offline setting, V-function is generally learned more reliably than the action-value function, which can provide better evaluation of an action indirectly.
Under this definition, an action $\bm{a}$ is considered advantageous if it leads to a next state with a higher expected value than the selective threshold defined by the $\kappa$-th quantile. A positive advantage indicates that the action is favored, while a negative advantage indicates that the action is penalized. The parameter $\kappa$ controls the level of conservatism: larger values lead to higher thresholds and encourage conservatism, while smaller values promote optimism by more readily rewarding unseen actions. Notably, all the components in Eq. (9) are learned solely from the offline dataset.
@@
Offline RL algorithms based on the standard Bellman backup suffer from action distribution shift during training. This shift arises because the target values in Eq. (10) use actions sampled from the learned policy $\pi_{\theta}$, while the Q-function is trained only on actions sampled from the behavior policy that produced the dataset ($(\bm{s},\bm{a})\in\mathcal{D}$). By augmenting the standard Bellman operator with the advantage term $A(\bm{a^{\prime}}|\bm{s^{\prime}})$, we can effectively mitigate estimation errors in Q-function at OOD actions.
Moreover, we can show that the advantage-based Bellman operator $\mathcal{T}_{A}^{\pi_{\theta}}$ is still contractive.
@@
$\mathcal{T}_{A}^{\pi_{\theta}}$ is $\gamma$-contractive with respect to the $L_{\infty}$ norm, which has a unique fixed point.
We denote the unique fixed point of Eq. (10) by $Q_{\pi_{\theta}}^{A}$, and the normal Q-function of $\pi_{\theta}$ by $Q_{\pi_{\theta}}$. The following proposition provides the bound of their difference.
@@
$\|Q_{\pi_{\theta}}^{A}-Q_{\pi_{\theta}}\|_{\infty}\leq 2\lambda R_{\rm max}(1-\gamma)^{-2}.$
Figure 2: Comparison of prior evaluation methods against our advantage-based evaluation. We visualize the Q-function for a fixed state. The thick blue solid line denotes the optimal value function. The solid line denotes the Q-function learned via advantage-based evaluation. The dashed line denotes the Q-function learned via conservative evaluation. The dotted line denotes the Q-function learned via standard Bellman backup evaluation.
We further illustrate the effectiveness of our advantage-based evaluation in Figure˜2, where different line styles correspond to different evaluation methods. In the offline RL setting, due to the prohibition against interacting with the environment, directly applying standard Bellman backup (dotted line in Figure˜2) results in an erroneous Q-function, which tends to overestimate the value of OOD actions and thus leads to an ineffective policy. Meanwhile, conservative evaluation (dashed line) indiscriminately penalizes all OOD actions, resulting in suppressed Q-values across these actions and limiting the learned policy to a sub-optimal policy near the support of the dataset. By contrast, our advantage-based evaluation (solid line) defines an advantage function that effectively modulates the Q-function obtained from Bellman backup, enabling the policy to discover optimal actions even beyond the support of the dataset. This phenomenon is further validated by empirical results on the PointMaze tasks, as shown in Section˜5.2.
@@ 4 Advantage-Based Diffusion Actor-Critic Algorithm
Building on the preceding analysis, we now introduce Advantage-based Diffusion Actor-Critic (ADAC).
Diffusion Policy. We model our policy as the reverse process of a conditional diffusion model (Wang et al., 2022):
@@
where the terminal sample $a^{0}$ is used as the action for RL evaluation. During training, we sample $(\bm{s},\bm{a})$ pairs from the offline dataset $\mathcal{D}$ and construct noisy samples $\bm{a^{i}}=\sqrt{\bar{\alpha}_{i}}\bm{a}+\sqrt{1-\bar{\alpha}_{i}}\bm{\epsilon}$ (Eq. (3)), where $\alpha_{i}=1-\beta_{i}$, $\bar{\alpha}_{i}=\prod_{j=1}^{i}\alpha_{j}$, and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. Following DDPM (Ho et al., 2020), we train the following noise prediction model $\bm{\epsilon}_{\theta}(\bm{a^{i}},\bm{s},i)$ to approximate the added noise, which determines the reverse process $p_{\theta}(\bm{a^{i-1}}\,|\,\bm{a^{i}},\bm{s})$:
$\mathcal{L}_{\rm BC}(\theta)=\mathbb{E}_{i\sim\mathcal{U},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}),(\bm{s},\bm{a})\sim\mathcal{D}}\left[\big\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{i}}\bm{a}+\sqrt{1-\bar{\alpha}_{i}}\bm{\epsilon},\bm{s},i)\big\|^{2}\right],$ (11)
where $\mathcal{U}$ denotes the uniform distribution over $\{1,\dots,T\}$. $\mathcal{L}_{\rm BC}(\theta)$ is a behavior cloning (BC) loss, and minimizing it enables the diffusion model to learn the behavior policy $\mu$. At inference time, an action $\bm{a^{0}}$ is generated by sampling $a^{T}\sim\mathcal{N}(\bm{0},\bm{I})$ and iteratively applying the learned reverse process.
@@
7:  Update $\mathcal{L}_{\rm CRITIC}(\phi)$ according to Eq. (14) // Critic Update
8:  Update $\mathcal{L}_{\rm ACTOR}(\theta)$ according to Eq. (15)// Actor Update
@@
Advantage. In our practical implementation, the value function is parameterized by a neural network with parameters $\varphi$ and trained by minimizing the following expectile regression loss:
$\mathcal{L}_{\rm VALUE}(\varphi)=\mathbb{E}_{(\bm{s},\bm{a},r,\bm{s^{\prime}})\sim\mathcal{D}}\big[L_{2}^{\tau}(r+\gamma V_{\varphi}(\bm{s^{\prime}})-V_{\varphi}(\bm{s}))\big].$ (12)
@@
$\mathcal{L}_{\rm MODEL}(\psi)=\mathbb{E}_{(\bm{s},\bm{a},\bm{s^{\prime}})\sim\mathcal{D}}\big[\|P_{\psi}(\bm{s},\bm{a})-\bm{s^{\prime}}\|^{2}\big].$ (13)
Therefore, the behavior policy $\mu$, the value function $V$, and the transition dynamics $P$ can be learned by minimizing Eq. (11), Eq. (12), and Eq. (13), respectively. The advantage function $A(\bm{a}\,|\,\bm{s})$ is then computed as defined in Eq. (9). All components are trained jointly using only offline data and then kept fixed during subsequent actor–critic updates, making this stage computationally inexpensive.
@@
$\mathcal{L}_{\rm CRITIC}(\phi)=\mathbb{E}_{(\bm{s},\bm{a},\bm{s^{\prime}})\sim\mathcal{D},\bm{a^{\prime}}\sim\pi_{\theta}(\cdot\mid\bm{s^{\prime}})}\left[\left(r(\bm{s},\bm{a})+\gamma\big(Q_{\phi}(\bm{s^{\prime}},\bm{a^{\prime}})+\lambda A(\bm{a^{\prime}}\mid\bm{s^{\prime}})\big)-Q_{\phi}(\bm{s},\bm{a})\right)^{2}\right].$ (14)
@@
$\mathcal{L}_{\rm ACTOR}(\theta)=\mathcal{L}_{\rm BC}(\theta)-\alpha\mathbb{E}_{\bm{s}\sim\mathcal{D},\bm{a}\sim\pi_{\theta}}\left[Q_{\phi}(\bm{s},\bm{a})\right].$ (15)
We summarize our implementation in Algorithm˜1. A central feature of our method is the incorporation of $A(\bm{a}\,|\,\bm{s})$, which distinguishes it from prior approaches such as DQL (Wang et al., 2022) that rely solely on the standard Bellman backup.
@@ 5 Experiments
In this section, we begin by evaluating our method on the widely recognized D4RL benchmark (Fu et al., 2020). We then design a dedicated experiment on the D4RL task PointMaze to better visualize ADAC’s ability to identify beneficial OOD actions. Finally, we perform an ablation study to dissect the contribution of key components in our method.
@@
Baseline. We consider a diverse array of baseline methods that exhibits strong results in each domain of tasks. For policy regularization-based method, we compare with the classic BC, TD3+BC (Fujimoto & Gu, 2021), BEAR (Kumar et al., 2019), BRAC (Wu et al., 2019), BCQ (Fujimoto et al., 2019), AWR (Peng et al., 2019), O-RL (Brandfonbrener et al., 2021), and DQL (Wang et al., 2022). For pessimistic value function-based approach, we include CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2021), and REM (Agarwal et al., 2020). For model-based offline RL, we choose MoRel (Kidambi et al., 2020). For the classic online method, we include SAC (Haarnoja et al., 2018). For conditional sequence modeling approaches, we include DT (Chen et al., 2021), Diffuser (Janner et al., 2022), and DD (Ajay et al., 2022). We report the performance of baseline methods either from the best results published in their respective papers or from (Wang et al., 2022).
@@
[…]
@@ 5.1 Benchmark Results
Our method is evaluated on four task domains, with results summarized in Table˜1. We also provide domain-specific analysis to highlight key performance characteristics.
@@
Results for AntMaze Tasks. We follow the D4RL evaluation protocol with normalized scores, where $100$ corresponds to an expert policy (Fu et al., 2020). AntMaze is particularly challenging due to sparse rewards and the prevalence of suboptimal trajectories, which makes controlled exploration of OOD actions crucial. Under these conditions, ADAC delivers over 15% improvements across all baselines and task variants. We further observe smoother learning curves and fewer training collapses compared to DQL (Figure˜3), suggesting that advantage-guided updates help the agent discover and reliably exploit the small set of reward-yielding behaviors despite limited feedback.
Results for Gym Tasks. In the dense-reward gym mujoco locomotion suite, ADAC provides consistent gains on top of already strong baselines. HalfCheetah is the most challenging family in this suite and exhibits the largest relative improvement at roughly 10%, while Hopper and Walker2d also benefit. Because scores are normalized (expert =100=100, from a converged online policy), averages near or above $90$ indicate behavior approaching expert quality; ADAC pushes more tasks into this regime. Qualitatively, we find that the advantage signal helps curb over-optimistic updates on hard-to-model transitions while preserving high-value in-distribution behaviors, yielding both higher final returns and more stable training.
Results for Adroit and Kitchen Tasks. Adroit involves dexterous hand manipulation and is particularly prone to extrapolation error because human demonstrations cover a narrow region of the state–action space. While both DQL and ADAC use policy regularization, adding the advantage further curbs OOD over-optimism, yielding about 20% improvement over strong baselines. Kitchen uses a Franka arm to execute long-horizon, compositional goals, and we observe consistent gains there as well.
Together, these results indicate that advantage-guided updates translate beyond locomotion and maze navigation, improving reliability in settings that stress dexterous manipulation and sequential goal completion.